OpenStudy (anonymous):

Rewrite the expression as an equivalent expression that does not contain powers of trigonometric functions greater than 1 8 cos^2 x

OpenStudy (anonymous):

Use the trig identity: \[\cos ^{2}u = \frac{ 1+\cos(2u) }{ 2 }\]